# Enhancing Consistency and Mitigating Bias: A Data Replay Approach for Incremental Learning

### v1

@@ Abstract
Deep learning systems are prone to catastrophic forgetting when learning from a sequence of tasks, where old data from experienced tasks is unavailable when learning from a new task. To mitigate the problem, a line of methods propose to replay the data of experienced tasks when learning new tasks. These methods usually adopt an extra memory to store the data for replay. However, it is not expected in practice considering the memory constraint or data privacy issue. As a replacement, data-free data replay methods are proposed by inverting samples from the classification model. Though achieving good results, these methods still suffer from the inconsistency of the inverted and real training data, which is neglected in the inversion stage in recent works. To that effect, we propose to measure the data consistency quantitatively by some simplification and assumptions. Using the measurement, we analyze existing techniques for inverting samples and get some insightful information that inspires a novel loss function to reduce the inconsistency. Specifically, the loss minimizes the KL divergence of the distributions of inverted and real data under the tied multivariate Gaussian assumption, which is easy to implement in continual learning. In addition, we observe that the norms of old class weights turn to decrease continually as learning progresses. We thus analyze the underlying reasons and propose a simple regularization term to balance the class weights so that the samples of old classes are more distinguishable. To conclude, we propose the Consistency enhanced data replay with debiased classifier for Class Incremental Learning (CCIL). Extensive experiments on CIFAR-100, Tiny-ImageNet, and ImageNet100 show consistently improved performance of CCIL compared to previous approaches.
Index Terms:
@@
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
@@
To avoid catastrophic forgetting, some researchers proposed revising the old knowledge by replaying the data of experienced tasks when learning the new task. Some works achieved this goal by keeping an extra data memory with a fixed size and designing specific strategies to manage and learn from the limited data in the memory [6, 7, 8]. Different from them, other methods tried to incrementally train a generator to generate old samples [9, 10, 11] and learn from the generated data, which is not constrained by the number of samples but by the generation quality. What is more, they still needed memory for saving the generator in the whole incremental learning process. Regarding memory constraints and latent privacy issues in practice, both the two lines of methods are not that applicable.
@@
To fulfill these visions, there are some works [12, 13] that aim at performing class incremental learning in a data-free manner, called data-free class incremental learning (DFCIL). That is, neither extra data memory nor an extra generation model is available in the data replay stage. These works mainly build upon the idea of inversion-based data-free knowledge distillation [14, 15, 16] by inverting samples directly from the classification model. It is natural to introduce the technique into data-free CIL, where at each new learning phase, the trained model on experienced tasks is available but the previous data is inaccessible. As for these data replay-based methods, the memory constraint problem and latent privacy issues are sidestepped properly. But they still face the inversion quality problem, which is severe for the lack of data. Thus, existing methods made great efforts on how to distill from these inverted samples and achieve SOTA performance.
@@
Despite their success, the inversion quality problem further incurs data inconsistency between synthetic and real samples, which is left unsolved and ignored by existing methods. Different from them, we focus on the inversion stage and consider how to narrow the distribution gap, which is orthogonal to previous works. The schematic illustration of our motivation can be found in Figure 1. We argue that the efforts in the inversion stage will help the overall performance. Because in the ideal case, the distribution of synthetic data is the same as the real data and the CIL problem degenerates into an offline supervised learning problem, which is much easier to solve. To that effect, we propose two metrics to quantitatively measure the data consistency by some simplification and assumptions. With them, we take a closer look at the existing techniques and find that there is still room to further improve data consistency. It inspires a new loss term. Specifically, we model the data distribution in the representation space of the penultimate layer and align the estimated distributions of old real and synthetic samples accordingly. It is natural to implement in the incremental learning setting: we estimate the parameters of the distributions at the end of the old task and when the new task comes, we regularize the synthetic samples to have similar statistics as the estimated ones to enhance consistency.
@@
On the other hand, we observe an unexpected phenomenon during incremental learning, where the norm of weight vectors belonging to old classes is lower than that belonging to new classes. It is similar to the observation in [17], but the underlying reasons are different. In [17], the class imbalance between replay data and new data is contributed to the biased class vectors. However, in inversion-based methods, the different learning strategies and different properties of synthetic data and real data take the major responsibility. Unlike [17] where a post-processing refinement of class vectors is introduced, we propose to add a regularization term to align the unbiased class weights. In this way, the network and weight vectors are optimized together in an unbiased and matching manner, which is more friendly to the inversion stage.
@@
  • We measure the data consistency of synthetic and real old data with some simplification. Using the measurement, we analyze existing inversion losses and draw insightful information, which inspires us to enhance the data consistency with a novel loss term.
@@
The rest of the paper is unfolded as follows. Section II introduces related works of our method, including works for class incremental learning and consistency measurement. Section IV introduces the preliminaries of our work, presents the overall framework, and depicts consistency enhanced replay and the debiased classifier, and Section V shows the comparison results of our proposed method with others, ablation study, and parameter analysis. Section VI concludes this study.
II Related Work
II-A Class Incremental Learning
@@
There is a rich literature in class incremental learning and methods of different categories try to solve the catastrophic forgetting problem of CIL from different views. Some works [18, 19, 20, 21, 22, 23] propose to incrementally expand the network to adapt to classes of different tasks and the modules can perform as a whole to classify all the seen classes during the testing stage. One other effective strategy to alleviate catastrophic forgetting is to regularize the optimization of the deep model when learning new tasks. Lots of works [24, 25, 26, 27] in this line are proposed based on different understandings of how to maintain the important information of past tasks. What is more, some research improves the performance of CIL with the help of other extra resources like extra unlabeled data [28, 29], pretrained models [20], or prompts [30].
@@
Methods [31, 32, 33, 6, 7, 34] mostly related to ours are based on knowledge distillation [35], where data representing past tasks are available to replay. To name some, iCaRL [6] performed logits distillation between the old model and the new model on both the data stored in the memory and data of the new task, UCIR [33] proposed to distill the features between the old model and the new model, PODNet [7] alternatively chose to distill the pooled feature maps in multiple layers, and INP-LPP [34] preserved both the global property of the whole set and the ranking of the label with instance neighborhood-preserving loss and label priority-preserving loss during distillation. Instead of storing data in memory, some works [10] additionally keep a generative model in memory to generate enough samples for distillation. Despite their effectiveness and popularity, all these methods need a relatively big memory budget to perform well.
II-B Data-free Replay-based Class Incremental Learning
@@
Data-free class incremental learning refers to CIL without storing either generative models or training data from past tasks. To tackle this, LwF [36] substituted data in the new task for old encountered data for distillation. Owing to the distribution difference between new task data and old task data, the performance of LwF is relatively limited. On top of that, [37] proposed to only perform distillation on a subset of new data samples selected by normalized cosine scores between samples and saved prototypes. To close the gap between data for distillation and data encountered, [14] inverted the data for distillation from the classification model independently. However, this method failed when applied to the standard benchmark of class incremental learning. The underlying reason is the gap between the synthetic samples and real samples. To facilitate the learning, [12] designed the modified cross-entropy training and importance-weighted feature distillation. The following work [13] split the representation learning stage and classification refinement stage. During representation learning, [13] adopted a relation-guided distillation loss to alleviate the conflict between plasticity and stability. What is more, [38] built upon the inversion stage proposed in [15] and proposed an additional strategy to guarantee the quality of inverted samples by virtue of the memorized confusion matrix.
II-C Data Consistency Measurement
@@
GAN [39] loss is a widely-used and effective loss to generate samples that are indistinguishable from target samples. However, it is not suitable in our setting where no target data is available. Domain alignment methods [40, 41, 42] in domain generalization are also related to the measurement of data consistency. But unlike generating consistent data in our setting, domain alignment methods align data of different domains in the feature space with the help of additional models and/or data. The setting most relative to our work is the inference-time OOD detection [43, 44, 45, 46], where a pretrained classification model is frozen and training data is accessible (just like the end of the task in incremental learning). Our work builds on the assumptions proposed in [43] and proposes to align the statistical parameters in the feature space to enhance data consistency.
III Preliminaries
III-A Problem Settings
@@
The purpose of class incremental learning is to learn a model that can tell the classes of all encountered training samples that come as a sequence of tasks. We formulate it as follows. Suppose $\mathcal{T}_{0},\mathcal{T}_{1},\ldots,\mathcal{T}_{N-1}$ are the $N$ tasks to be learned sequentially. The label set $\mathcal{C}_{i}$ of $\mathcal{T}_{i}$ meets the following condition: $\forall i,j,i\neq j,\mathcal{C}_{i}\cap\mathcal{C}_{j}=\varnothing$, which indicates that two samples from different tasks can not belong to the same class. After learning $\mathcal{T}_{0},\mathcal{T}_{1},\ldots,\mathcal{T}_{N-1}$ (denoted as $\mathcal{T}_{0:N-1}$), the classification model is supposed to classify samples from all seen classes $\mathcal{C}_{0}\cup\mathcal{C}_{1}\cup\ldots\cup\mathcal{C}_{N-1}$ (denoted as $\mathcal{C}_{0:N-1}$). On top of that, at task $\mathcal{T}_{i}$ in our setting, samples belonging to classes in $\mathcal{C}_{i}$ are sufficient while no samples belonging to past tasks $\mathcal{T}_{0:i-1}$ are accessible. The classification model to be trained is denoted as $f(\cdot)$ which is composed of a feature extractor $h(\cdot)$ which has the deep neural network architecture, and a classifier $g(\cdot)$ which is the final fully-connected layer, that meets: $f(\cdot)=g(h(\cdot))$.
III-B Baseline Approach
In this section, we introduce a baseline method. For the inversion stage, the objective is to invert input samples from a classification model. To that effect, we froze the model and optimize the pixels of inputs according to some losses on the features and outputs. We adopt DeepInversion [14] in this baseline just like other SOTA DFCIL methods [13, 12].
@@ III-B Baseline Approach
The first component is the cross-entropy loss based on the assumption that inverted samples can be classified correctly by the old model. Let us denote the inverted samples at $\mathcal{T}_{i}$ as $\hat{X}^{i-1}$, the old model as $f_{0:i-1}$. The cross-entropy loss is given as follows:
$\mathcal{L}_{ce}=\frac{1}{|\hat{X}^{i-1}|}\sum_{(\hat{x},\hat{y})\in\hat{X}^{i-1}}\ell_{ce}(f_{0:i-1}(\hat{x}),\hat{y},\tau),$ (1)
where $\tau$ is the temperature.
@@
Besides, the statistics alignment loss is proposed to align the Batch Normalization (BN) statistics. Each BN layer stores the running mean and variance of features, which can be directly utilized as the prior knowledge to regularize the inversion stage as follows:
@@
where $l$ indicates the $l$-th BN layer in the model, $\boldsymbol{\sigma_{l}},\boldsymbol{\mu_{l}}$ denotes the mean and variance of the $i$-th BN layer stored in the old model, $\boldsymbol{\hat{\sigma}_{l}},\boldsymbol{\hat{\mu}_{l}}$ denotes the mean and variance estimated on the inverted data, and $\mathcal{D}_{KL}$ denotes the KL divergence.
Following [12, 13], to improve the efficiency of the inversion, a generator $G(\cdot)$ with trainable parameters is introduced to generate input samples: $\hat{x}=G(n),\hat{y}=\mathop{\mathrm{argmax}}\limits_{k}f_{0:i-1}(\hat{x})_{k}$, where $n$ is the noise. By introducing the generator, parameters to be optimized change from pixels of inputs $\hat{X}^{i-1}$ to parameters of $G(\cdot)$. To guarantee class balance in the generated data, a class diversity loss is adopted in the baseline as follows:
$\mathcal{L}_{div}=H\left(\frac{1}{|\hat{X}^{i-1}|}\sum_{(\hat{x},\hat{y})\in\hat{X}^{i-1}}{\rm Softmax}\left(f_{0:i-1}(\hat{x})\right)\right),$ (3)
where $H$ denotes the information entropy. We remark that the generator is not necessary for the inversion stage and one can directly optimize $\hat{X}^{i-1}$ with a sacrifice to the inversion efficiency. Also, note that, unlike generative data replay [10], the generator here is inverted only from $f_{0:i-1}$ and can be immediately discarded after data inversion, which is viewed as data-free in CIL.
The overall loss for the inversion stage is:
@@
where $\lambda_{1},\lambda_{2}$ are hyperparameters controlling the scales of the losses.
As for the training stage, we adopt R-DFCIL [13] as our baseline because of its SOTA performance. In R-DFCIL, hard knowledge distillation (HKD) is proposed to distill the information of old tasks, local cross-entropy loss (LCE) is adopted for learning from samples of current tasks, and relational knowledge distillation loss builds a bridge for old and new models when training with new samples. We denote the baseline training loss as $\mathcal{L}_{T\_base}$. To be noticed, our method is orthogonal to previous works contributing to the distillation losses and can be combined with any of them (See details in Section V-B).
The details of HKD and LCE are provided here for the latter analysis. HKD is applied on synthetic samples to prevent model changes as follows:
$\mathcal{L}_{hkd}=\frac{1}{|\hat{X}^{i-1}|\times|\mathcal{C}_{0:i-1}|}\sum_{(\hat{x},\hat{y})\in\hat{X}^{i-1}}||f_{0:i-1}(\hat{x})-f_{0:i}(\hat{x})||_{1}.$ (5)
@@
Then, a cross-entropy loss is applied locally to the current task’s samples (denoted as $X^{i}$) as follows:
$\mathcal{L}_{lce}=\frac{1}{|X^{i}|}\sum_{(x,y)\in X^{i}}\ell_{ce}(f_{i}(x),y,\tau),$ (6)
where $f_{i}(\cdot)=g_{i}(h_{0:i}(\cdot))$, that apply Softmax on the logits of classes belonging to the current task.
IV Method
IV-A Overall Framework

### v2

@@ Abstract
Deep learning systems are prone to catastrophic forgetting when learning from a sequence of tasks, as old data from previous tasks is unavailable when learning a new task. To address this, some methods propose replaying data from previous tasks during new task learning, typically using extra memory to store replay data. However, it is not expected in practice due to memory constraints and data privacy issues. Instead, data-free replay methods invert samples from the classification model. While effective, these methods face inconsistencies between inverted and real training data, overlooked in recent works. To that effect, we propose to measure the data consistency quantitatively by some simplification and assumptions. Using this measurement, we gain insight to develop a novel loss function that reduces inconsistency. Specifically, the loss minimizes the KL divergence between distributions of inverted and real data under a tied multivariate Gaussian assumption, which is simple to implement in continual learning. Additionally, we observe that old class weight norms decrease continually as learning progresses. We analyze the reasons and propose a regularization term to balance class weights, making old class samples more distinguishable. To conclude, we introduce Consistency-enhanced data replay with a Debiased classifier for class incremental learning (CwD). Extensive experiments on CIFAR-100, Tiny-ImageNet, and ImageNet100 show consistently improved performance of CwD compared to previous approaches.
keywords:
@@
journal: Neural Networks
1 Introduction
@@
To avoid catastrophic forgetting, some researchers proposed revising the old knowledge by replaying the data of the experienced tasks when learning the new task. Some works achieved this goal by keeping an extra data memory with a fixed size and designing specific strategies to manage and learn from the limited data in the memory [6, 7, 8, 9, 10]. Unlike them, other methods tried to incrementally train a generator to generate old samples [11, 12, 13] and learn from the generated data, which is not constrained by the number of samples but by the quality of the generation. In addition, they still needed memory to save the generator in the whole incremental learning process. Regarding memory constraints and latent privacy issues in practice, both the two lines of methods are not that applicable.
@@
To overcome limitations, some works [14, 15] aim to perform class incremental learning in a data-free manner, called data-free class incremental learning (DFCIL). That is, neither additional data nor models are carried over across tasks. These works mainly build on the idea of inversion-based data-free knowledge distillation [16, 17, 18] by inverting samples directly from the classification model. It is natural to introduce the technique into data-free CIL, where at each new learning phase the trained model on experienced tasks is available but the previous data are inaccessible. As for these data-replay-based methods, the memory constraint problem and latent privacy issues are properly sidestepped. However, they still face an inversion quality problem which is severe because of the lack of data. Thus, existing methods have made great efforts on how to distill from these inverted samples and achieve SOTA performance.
@@
Despite their success, the inversion quality problem further causes data inconsistency between synthetic and real samples, which is left unsolved and ignored by existing methods. Unlike them, we focus on the inversion stage and consider how to narrow the distribution gap, which is orthogonal to previous works. The schematic illustration of our motivation can be found in Figure 1. We argue that efforts in the inversion stage will help the overall performance. In the ideal case, the distribution of synthetic data is the same as the real data and the CIL problem degenerates into an offline supervised learning problem, which is much easier to solve. To that effect, we propose to quantitatively measure the data consistency by some simplification and assumptions. Specifically, we model the data distribution in the representation space of the penultimate layer under a tied multivariate Gaussian assumption, which inspires a new loss term by aligning the estimated distributions of old real and synthetic samples accordingly. It is natural to implement in the incremental learning setting: we estimate the parameters of the distributions at the end of the old task and when the new task comes, we regularize the synthetic samples to have similar statistics as the estimated ones to enhance consistency.
@@
On the other hand, we observe an unexpected phenomenon during incremental learning, where the norm of weight vectors belonging to old classes is lower than that belonging to new classes. It is similar to the observation in [19], but the underlying reasons are different. In [19], the class imbalance between replay data and new data contributed to the biased class vectors. However, in inversion-based methods, the different learning strategies and different properties of synthetic data and real data take the major responsibility. Unlike [19] where a post-processing refinement of class vectors is introduced, we propose to add a regularization term to align the unbiased class weights. In this way, the network and weight vectors are optimized together in an unbiased and matching manner, which is more friendly to the inversion stage.
@@
  • 1.
    We quantitatively measure the data consistency of synthetic and real old data with some simplification, which inspires a novel loss term for enhancing the data consistency.
  • 2.
@@
The structure of the paper is as follows. In Section 2, we discuss related work, including class-incremental learning methods and consistency measurement. Section 3 outlines the problem setting and DFCIL baselines. Section 4 describes the proposed framework, focusing on consistency-enhanced replay and the debiased classifier. Section 5 presents experimental comparisons, ablation studies, and further analysis. In Section 6, we discuss the potential limitations of our approach, and finally, Section 7 summarizes our findings.
2 Related Work
2.1 Class Incremental Learning
@@
There is a rich body of literature on class incremental learning (CIL), with various methods addressing the catastrophic forgetting problem from different perspectives. Some works [20, 21, 22, 23, 24, 25, 26] proposed incrementally expanding the network to accommodate new classes from different tasks. These methods allow the modules to function collectively during the testing stage to classify all previously seen classes. Another effective strategy to alleviate catastrophic forgetting is to regularize the optimization of the deep model when learning new tasks. Works in this area aimed to preserve crucial information from past tasks [27, 28, 29, 30] or to imitate the oracle model [31]. Additionally, some research enhanced the performance of CIL by leveraging extra resources, including unlabeled data [32, 33], pretrained models [24], and auxiliary tasks [34].
@@
The methods [6, 35, 7, 8, 9, 36, 37] closely related to ours are based on knowledge distillation [38], where data representing past tasks are available for replay. For example, iCaRL [6] performed logits distillation between old and new models using both the stored data and the data from the new task. UCIR [7] distilled features between the old and new models, PODNet [9] chose to distill pooled feature maps across multiple layers, and Wu et al. [37] introduced hyper-feature aggregation and relaxed distillation to improve knowledge retention and flexibility of adaptation. Instead of storing data, some works [12] additionally kept a generative model in memory to produce enough samples for distillation. Despite their effectiveness and popularity, all of these methods require a relatively large memory budget to perform well.
2.2 Data-free Class Incremental Learning
@@
Data-free class incremental learning refers to CIL without storing either generative models or training data from past tasks. To address this, LwF [39] substituted data from the new task for previously encountered data for distillation. Due to the distribution differences between new task data and old task data, LwF’s performance was relatively limited. In this regard, Zhu et al. [40] proposed performing distillation only on a subset of new data samples selected by saved prototypes. Shi and Ye [41] further updated class prototypes while improving knowledge transfer between old and new classes. To close the gap between data used for distillation and encountered data, replay-based methods [16, 14, 15, 42, 17] directly inverted data for distillation from the classification model. [16] was the first attempt but failed on the standard CIL benchmark due to the gap between synthetic and real samples. To facilitate learning, [14] designed modified cross-entropy training and importance-weighted feature distillation. The subsequent work [15] split the representation learning and classification refinement stages. During representation learning, [15] adopted a relation-guided distillation loss to alleviate the conflict between plasticity and stability. Additionally, [42] built upon the inversion stage proposed in [17] and introduced an additional strategy to ensure the quality of inverted samples using a memorized confusion matrix. The major distinction between our work and these approaches is the introduction of an estimation stage after training. By estimating statistics of past tasks’ data, we enhance the consistency between inverted data and the original data from previous tasks. Furthermore, we propose a weight alignment regularization term to address the classifier bias caused by the learning strategy and inverted samples, which is first reported in the literature.
2.3 Data Consistency Measurement
@@
GAN [43] loss is a widely used and effective loss to generate samples that are indistinguishable from target samples. However, it is not suitable in our setting where no target data is available. Domain alignment methods [44, 45, 46] in domain generalization are also related to the measurement of data consistency. But unlike generating consistent data in our setting, domain alignment methods align data of different domains in the feature space with the help of additional models and/or data. The setting most relative to our work is the inference-time OOD detection [47, 48, 49, 50], where a pretrained classification model is frozen and training data is accessible (just like the end of the task in incremental learning). Our work builds on the assumptions proposed in [47] and proposes to align the statistical parameters in the feature space to enhance data consistency.
3 Preliminaries
3.1 Problem Settings
@@
The purpose of class incremental learning is to learn a model that can tell the classes of all encountered training samples that come as a sequence of tasks. We formulate it as follows. Suppose $\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{N}$ are the $N$ tasks to be learned sequentially. The label sets are disjoint among different tasks: $\forall i,j,i\neq j,\mathcal{C}_{i}\cap\mathcal{C}_{j}=\varnothing$. After learning $\mathcal{T}_{1},\mathcal{T}_{2},\ldots,\mathcal{T}_{N}$ (denoted as $\mathcal{T}_{1:N}$), the classification model is supposed to classify samples from all seen classes $\mathcal{C}_{1}\cup\mathcal{C}_{2}\cup\ldots\cup\mathcal{C}_{N}$ (denoted as $\mathcal{C}_{1:N}$). On top of that, at task $\mathcal{T}_{i}$ in our setting, samples belonging to classes in $\mathcal{C}_{i}$ are sufficient while no samples belonging to past tasks $\mathcal{T}_{1:i-1}$ are accessible. The classification model, denoted as $f(\cdot)$, consists of a feature extractor $h(\cdot)$, which follows a deep neural network architecture, and a classifier $g(\cdot)$, which represents the final fully connected layer, such that $f(\cdot)=g(h(\cdot))$.
3.2 Inversion-based DFCIL Approaches
In this section, we describe the pipeline of inversion-based DFCIL approaches and the specific baseline on which our work is based. Typically, inversion-based methods consist of two main stages: the inversion stage and the training stage. In the DFCIL setting, due to the lack of previous data, the inversion-based method begins with the inversion stage, which directly optimizes samples or trains a generator based only on the old classification model. These optimized samples or the trained generator act as a surrogate for the real old data. Given that the classification model is traditionally designed to classify samples, this inverse data generation process based on it is termed inversion, and samples created in this way are referred to as inverted samples. After the inversion stage, the new task data and inverted data (either offline optimized samples in the inversion stage or online generated data by the trained generator) are used to train the new classification model for all encountered tasks, which forms the training stage.
In chronological order, we briefly introduce representative inversion-based works from the literature. DeepInversion [16] was the initial attempt, where samples were directly optimized (treating pixels as parameters) during the inversion stage, and the standard cross-entropy (CE) loss was adopted for these inverted samples. ABD [14] extended DeepInversion by making two key modifications: training a generator instead of samples during the inversion stage for efficiency and adopting importance-weighted feature distillation to overcome the problem caused by the gap between the inverted and real samples. R-DFCIL [15] retained the same inversion stage as ABD but introduced a relation-guided distillation loss during the training stage to better balance plasticity and stability.
In this paper, we combine the inversion stage proposed in ABD and the training stage proposed in R-DFCIL as our baseline method due to its SOTA performance. To be noticed, the inversion stage of ABD used the same objective function as DeepInversion with different optimizable objects. As a result, the inversion stage is still called DeepInversion.
@@ 3.2 Inversion-based DFCIL Approaches
The first component of DeepInversion is the cross-entropy loss based on the assumption that inverted samples can be classified correctly by the old model. Let us denote the generator to be trained as $G(\cdot)$ and the old model as $f_{i-1}(\cdot)=g_{1:i-1}(h_{i-1}(\cdot))$. Each online inverted sample is generated as follows: $\hat{x}=G(n),\hat{y}=\mathop{\mathrm{argmax}}\limits_{k}f_{i-1}(\hat{x})_{k}$, where $n$ follows low-dimensional Gaussian distribution $\mathcal{N}$. Then the cross-entropy loss is given as follows:
$\mathcal{L}_{ce}=\mathop{\mathbb{E}}\limits_{n\sim\mathcal{N}}[\ell_{ce}(f_{i-1}(G(n)),\mathop{\mathrm{argmax}}\limits_{k}f_{i-1}(G(n))_{k},\tau)],$ (1)
@@
where $\tau$ is the temperature. Besides, the statistics alignment loss is proposed to align the Batch Normalization (BN) statistics. Each BN layer stores the running mean and variance of features, which can be directly utilized as the prior knowledge to regularize the inversion stage as follows:
@@
where $l$ indicates the $l$-th BN layer in the model, $\boldsymbol{\sigma_{l}},\boldsymbol{\mu_{l}}$ denotes the mean and variance of the $i$-th BN layer stored in the old model, $\boldsymbol{\hat{\sigma}_{l}},\boldsymbol{\hat{\mu}_{l}}$ denotes the mean and variance estimated on the inverted data, and $\mathcal{D}_{KL}$ denotes the KL divergence. Furthermore, to ensure class balance in the generated data, a class diversity loss is adopted in the baseline as follows:
$\mathcal{L}_{div}=H\left(\mathop{\mathbb{E}}\limits_{n\sim\mathcal{N}}\left[{\rm Softmax}(f_{i-1}(G(n)))\right]\right),$ (3)
where $H$ denotes the information entropy. The overall loss for training $G(\cdot)$ in the inversion stage is:
@@
where $\lambda_{1},\lambda_{2}$ are hyperparameters controlling the scales of the losses. It is worth noting that unlike generative data replay [12], the generator used here is inverted from the old classification model $f_{i-1}$ in the inversion stage of the $i$-th task and can be immediately discarded after the training stage of the $i$-th task, as it is not stored for future use. In other words, a generator is randomly initialized and trained every time a new task comes.
During the training stage, we adopted R-DFCIL [15] as our baseline. In R-DFCIL, hard knowledge distillation (HKD) was used to distill information from old tasks, local cross-entropy loss (LCE) was employed for learning from current task samples, and relational knowledge distillation (RKD) loss created a connection between old and new models when training with new samples. Specifically, HKD is applied on synthetic samples to prevent model changes as follows:
$\mathcal{L}_{hkd}=\frac{1}{|\mathcal{C}_{1:i-1}|}\mathop{\mathbb{E}}\limits_{n\sim\mathcal{N}}||f_{i-1}(G(n))-[f_{i}(G(n))]_{1:|\mathcal{C}_{1:i-1}|}||_{1},$ (5)
@@
where $[f_{i}(G(n))]_{1:|\mathcal{C}_{1:i-1}|}$ denotes the sub-vector of $f_{i}(G(n))$ with dimensions from 1 to $|\mathcal{C}_{1:i-1}|$. Then, a cross-entropy loss is applied locally to the current task’s samples (denoted as $X^{i}$) as follows:
$\mathcal{L}_{lce}=\frac{1}{|X^{i}|}\sum_{(x,y)\in X^{i}}\ell_{ce}(g_{i}(h_{i}(x)),y,\tau),$ (6)
where $g_{i}(h_{i}(x))$ means applying Softmax on the logits of classes belonging to the current $i$-th task. We omit the introduction of the RKD loss for its complexity and its limited relevance to the focus of our work. We denote the baseline training loss $\mathcal{L}_{T\_base}$, which consists of these three terms, as follows:
$\mathcal{L}_{T\_base}=\lambda_{3}\mathcal{L}_{hkd}+\lambda_{4}\mathcal{L}_{lce}+\lambda_{5}\mathcal{L}_{rkd},$ (7)
where $\lambda_{3},\lambda_{4},\lambda_{5}$ are corresponding coefficients. Notably, our contributions are orthogonal to previous works’ that relate to the distillation loss in the training stage and can be combined with any of them (see details in Section 5.2). When the choice of the baseline method changes, $\mathcal{L}_{T\_base}$ will change accordingly.
4 Method
4.1 Overall Framework